You have a deck of 52 playing cards (check online for the full composition of a 52 card deck if you are unsure).

(i) How many different 8 card hands can be dealt?

(ii) What is the probability that a hand of 8 dealt randomly contains (exactly) 2 aces?

(iii) What is the probability that a hand of 7 dealt randomly will have 7 cards of the same suit? 


1
Expert's answer
2021-12-13T14:16:03-0500

"i)"

Using combinations, we can find the number of 8 card hands that can be dealt as follows,

number of ways ="\binom{n}{x}={n!\over(x!(n-x)!)}"

From the question, "n=52, x=8"

We now have,

"\binom{52}{8}={52!\over(8!\times44!)}=752538194"

Therefore, there are 752538194-8card hands that can be dealt


"ii)"

To get exactly 2 aces, we need to choose 2 out of 4 aces and 6 of the remaining 48 cards. The number of ways to do this is

"\binom{4}{2}\times\binom{48}{6}={4!\over(2!*2!)}*{48!\over(6!*42!)}=6\times12271512=73629072"

The probability is, "{73629072\over752538194}=0.0978(4dp)"

Therefore, probability of getting exactly 2 aces is 0.0978.


"iii)"

Since there are 13 cards in a suit, we choose 7 cards from the 13 cards using combinations as follows, "\binom{13}{7}={13!\over(7!*6!)}=1716"

There are 4 suits but we only want 1suit. Therefore, the number of ways for choosing 1 out of a total of 4 suits is, "\binom{4}{1}={4!\over(1!\times3!)}=4"

Therefore, total number of ways is 1716*4=6864

Hence the probability is, "{6864\over752538194}=9.121132e{-6}"

Therefore, the probability that a hand of 7 dealt randomly will have 7 cards of the same

suit is 9.121132e-6.